OpenStudy (anonymous):

solve this system show steps please so i can learn how to do this problem y=-x^2+5x+1 y=2x-3

OpenStudy (anonymous):

-x^2 + 5x + 1 = 2x - 3 //both are equal to y, so they're equal to each other -x^2 + 3x + 4 = 0 //move like terms to the left end -(x^2 - 3x - 4) = 0 //factor out -1 -(x+1)(x-4) = 0 //breaking it down into simpler terms x + 1 = 0 OR x -4 = 0 x = -1 or x = 4